Expected Returns 0.17 0.11 0.30 Standard Deviation 0.12 0.05 Firm As common stock Firm Bs common stock Correlation coeffici(Computing the standard deviation for a portfolio of two risky investments) Mary Guilott recently graduated from college and
Answer #1

a. Expected Return of Portfolio =10%*0.17+90%*0.11 =11.60%
Standard Deviation of Portfolio =((10%*0.12)^2+((90%*0.05)^2+2*10%*90%*0.12*0.05*0.30)^0.5 =4.99%

b. Expected Return of Portfolio =90%*0.17+10%*0.11 =16.40%
Standard Deviation of Portfolio =((90%*0.12)^2+((10%*0.05)^2+2*90%*10%*0.12*0.05*0.30)^0.5 =10.96%

c. Expected Return of Portfolio =10%*0.17+90%*0.11 =11.60%
Standard Deviation of Portfolio =((10%*0.12)^2+((90%*0.05)^2-2*10%*90%*0.12*0.05*0.30)^0.5 =4.30%

d. Expected Return of Portfolio =90%*0.17+10%*0.11 =16.40%
Standard Deviation of Portfolio =((90%*0.12)^2+((10%*0.05)^2-2*90%*10%*0.12*0.05*0.30)^0.5 =10.66%